radomes.openskiesproject.org
Juliegotmuddled18
Juliegotmuddled18 Juliegotmuddled18
  • 26-03-2015
  • Mathematics
Answered

If a probability of a win is 0.5 and the probability of a draw is 0.3, what is the probability of a loss?

Verified Answer :

BaekEmmaSakeLili
BaekEmmaSakeLili BaekEmmaSakeLili
  • 26-03-2015

The total has to add up to 1 in terms of probability,  so the decimal has to be 0.2.

0.5 + 0.3 + 0.2 = 1

Answer Link
Bryancohen
Bryancohen Bryancohen
  • 11-07-2019

Answer: 0.2

Step-by-step explanation: Adding the probability of a win and a draw, you would get 0.8.

0.5+0.3=0.8

The probability of a loss is going to be the rest. Subtract 0.8 from 1 to get the probability.

1-0.8=0.2

The probability of a loss is 0.2.

Answer Link

Related Questions

A living things place to live in an ecosystem
What can the writer do to improve the format and structure of this letter of complaint The format and structure of the letter is correct as is The senders name
All adult men in the Massachusetts Bay Colony who were members of the church were given the right to vote True False
Social impacts of urbanization
Which description best fits the Neolithic Revolution 1 moving from urban centers to rural centers 2 using petrochemical fertilizers and pesticides to increase p
Which of the following is an example of a declarative sentence
What Was Kristallnacht
What does the alt code 158 mean
The treatment based on the theory of pressure on blood vessels and nerves is a psychotherapy b chiropody c osteomyelitis d osteopathy
How is a solute different from a solvent in a solution